OpenStudy (anonymous):

Soybean milk is 16% protein and corn meal is 9% protein. How many pounds of each should be mixed together to get a 350-lb mixture that is 12% protein. Soybean milk is 16% protein and corn meal is 9% protein. How many pounds of each should be mixed together to get a 350-lb mixture that is 12% protein. @Mathematics

OpenStudy (anonymous):

let x=16% protein and y=9% protein 16%x+ 9%y=12%(350 lb) 0.16x+.09y=0.12(350) 1st equation x+y=350 2nd equation solving for x and y by substitution or elimination method x= 150 lb and y=200 lb

OpenStudy (anonymous):

Thank you... I couldn't figure out the equations :D I suck at translating word problems